OpenStudy (anonymous):

F(x)=4x^2+16x+15 find the x and y intercepts

OpenStudy (anonymous):

To find the y intercept you put 0 in to the x values. y = 4(0^2) + 16(0) + 15 y = 4(0) + 0 + 15 y = 0 + 0 + 15 y = 15 y-intercept (0,15)

OpenStudy (anonymous):

To find the x-intercepts you put y = 0 and factor. 0 = 4x^2 + 16x + 15 0 = (2x + 5)(2x + 3) 2x + 5 = 0 or 2x + 3 = 0 2x = -5 2x = -3 x = -5/2 x = -3/2 x-intercepts are (-2.5,0) and (-1.5,0)

OpenStudy (anonymous):

put 0 in to the x values